Nasir wants to accumulate $1500000 after 10 years. If investment earns interest at the rate of 8 percent per year compounded quarterly, then the amount he must be deposited at the end of each year is 

None of the above


 

$21,900

$22,354

$24,840

$23,220


1
Expert's answer
2021-02-02T04:54:49-0500

find the formula:

"FV=X\times\frac{(1+r)^n-1}{r}"


FV= 1 500 000

n=40, "4\times10=40"

"r=\frac{8.5}{4}=2.125"

"1 500 000=X\times\frac{(1+0.02125)^{40}-1}{0.02125}"


"1 500 000=X\times 62.0660734"

x=24 166.26


None of the above